OpenStudy (anonymous):

You put $600 into a savings account paying 8% annual interest compounded monthly. After one year you take out all the money and put it into a 9 month CD paying 7.5% annual(simple) interest. How much is your investment worth when the CD matures?

OpenStudy (anonymous):

$915.13

OpenStudy (anonymous):

can u quickly explain how u get it....did u use formula A=P (1+i) ^n

OpenStudy (anonymous):

i do and it wont let me...??

OpenStudy (anonymous):

A=600 A(1+(.08/12))^12= 866.40= B B(1+(.075*.75)) =915.14

OpenStudy (anonymous):

thank u sooo much =)